# CompdVision: Combining Near-Field 3D Visual and Tactile Sensing Using a Compact Compound-Eye Imaging System

### v2

@@ Abstract
As automation technologies advance, the need for compact and multi-modal sensors in robotic applications is growing. To address this demand, we introduce CompdVision, a novel sensor that employs a compound-eye imaging system to combine near-field 3D visual and tactile sensing within a compact form factor. CompdVision utilizes two types of vision units to address diverse sensing needs, eliminating the need for complex modality conversion. Stereo units with far-focus lenses can see through the transparent elastomer for depth estimation beyond the contact surface. Simultaneously, tactile units with near-focus lenses track the movement of markers embedded in the elastomer to obtain contact deformation. Experimental results validate the sensor’s superior performance in 3D visual and tactile sensing, proving its capability for reliable external object depth estimation and precise measurement of tangential and normal contact forces. The dual modalities and compact design make the sensor a versatile tool for robotic manipulation.
@@ I Introduction
Vision
Tactile
GelSight[5]
No
Force + Shape
30fps
35$\times$35$\times$60
OmniTact[9]
No
Shape
30fps
$\varnothing$30$\times$33
Zhang et al. [10]
No
Force
30fps
14.5$\times$14.5$\times$5
FingerVision[11]
RGB, limited
Force
Simultaneous sensing
30fps
47$\times$40$\times$30
Finger-STS[14]
RGB
Marker Flow
LEDs
NA
50$\times$50
SpecTac[12]
RGB
Force
LEDs
0.2s visual-tactile cycle
NA
StereoTac[15]
RGBD
Shape
LEDs
4Hz LEDs alternating rate
NA
CompdVision (Ours)
RGBD
Force
Simultaneous sensing
25fps
22$\times$14$\times$14
@@ V-B Force Measurement Calibration
After collecting data, we randomly split the dataset, allocating 70% for training and 30% for testing. Our data processing employed a Convolutional Neural Network (CNN) model with multiple convolutional and fully connected layers. The model incorporates five convolutional layers and batch normalization for stable learning. Enhanced by max pooling and dropout for regularization, the architecture concludes with fully connected layers, leading to three output neurons for the forces $[F_{x},F_{y},F_{z}]$, where $F_{x}$ and $F_{y}$ denote tangential forces, and $F_{z}$ denotes the normal force. The evaluation results using the testing dataset, as shown in Table III, indicate the RMSE for the force measurement on all three axes. The experiments and results underscore our sensor’s robust capability for accurate contact force measurement.

### v3

@@ Abstract
As automation technologies advance, the need for compact and multi-modal sensors in robotic applications is growing. To address this demand, we introduce CompdVision, a novel sensor that employs a compound-eye imaging system to combine near-field 3D visual and tactile sensing within a compact form factor. CompdVision utilizes two types of vision units to address diverse sensing needs, eliminating the need for complex modality conversion. Stereo units with far-focus lenses can see through the transparent elastomer for depth estimation beyond the contact surface. Simultaneously, tactile units with near-focus lenses track the movement of markers embedded in the elastomer to obtain contact deformation. Experimental results validate the sensor’s superior performance in 3D visual and tactile sensing, proving its capability for reliable external object depth estimation and precise measurement of tangential and normal contact forces. The dual modalities and compact design make the sensor a versatile tool for robotic manipulation. Videos are available at hkust-orisys.github.io/CompdVision.
@@ I Introduction
Vision Tactile
GelSight[5] No Force + Shape 30fps 35$\times$35$\times$60
OmniTact[9] No Shape 30fps $\varnothing$30$\times$33
Zhang et al. [10] No Force 30fps 14.5$\times$14.5$\times$5
FingerVision[11] RGB, limited Force Simultaneous sensing 30fps 47$\times$40$\times$30
Finger-STS[14] RGB Marker Flow LEDs NA 50$\times$50
SpecTac[12] RGB Force LEDs 0.2s visual-tactile cycle NA
StereoTac[15] RGBD Shape LEDs 4Hz LEDs alternating rate NA
CompdVision (Ours) RGBD Force Simultaneous sensing 25fps 22$\times$14$\times$14
@@ V-B Force Measurement Calibration
After collecting data, we randomly split the dataset, allocating 70% for training and 30% for testing. Our data processing employed a Convolutional Neural Network (CNN) model with multiple convolutional and fully connected layers. The model incorporates five convolutional layers and batch normalization for stable learning. Enhanced by max pooling and dropout for regularization, the architecture concludes with fully connected layers, leading to three output neurons for the forces $[F_{x},F_{y},F_{z}]$, where $F_{x}$ and $F_{y}$ denote tangential forces, and $F_{z}$ denotes the normal force. The evaluation results using the testing dataset, as shown in Table III, indicate the RMSE and the coefficient of determination ($R^{2}$) for the force measurement on all three axes. The experiments and results underscore our sensor’s robust capability for contact force measurement.